# Radio Observations as an Extrasolar Planet Discovery and Characterization: Interior Structure and Habitability

### v1

@@ Abstract
Detection of radio emission from Jupiter was identified quickly as being due to its planetary-scale magnetic field. Subsequent spacecraft investigations have revealed that many of the planets, and even some moons, either have or have had large-scale magnetic fields. In the case of the Earth, Jupiter, Saturn, Uranus, and Neptune, the their magnetic fields are generated by dynamo processes within these planets, and an interaction between the solar wind and their magnetic fields generates intense radio emission via the electron cyclotron maser instability. In the case of Jupiter, its magnetic field interacts with the moon Io to result in radio emission as well. Extrasolar planets reasonably may be expected to generate large-scale magnetic fields and to sustain an electron cyclotron maser instability. Not only may these radio emissions be a means for discovering extrasolar planets, because magnetic fields are tied to the properties of planetary interiors, radio emissions may be a remote sensing means of constraining extrasolar planetary properties that will be otherwise difficult to access. In the case of terrestrial planets, the presence or absence of a magnetic field may be an indicator for habitability. Since the first edition of the Handbook, there have been a number of advances, albeit there remain no unambigous detection of radio emission from extrasolar planets. New ground-based telescopes and new possibilities for space-based telescopes provide promise for the future.
@@ 0.1 Introduction
The focus of this chapter is on electron cyclotron maser instability (ECMI) emissions resulting from a stellar wind-planetary magnetosphere interaction. While planets, or at least Jupiter, can generate synchrotron emission that appears at decimeter and centimeter wavelengths, it is likely to remain well beyond detection capabilities in the near future. As an example, the centimeter-wavelength flux density of Jupiter is of order a few Janskys (e.g., Klein et al., 1989; de Pater et al., 2003). Scaled to a distance of even a few parsecs, the resulting flux density would be sub-nanoJanskyy, which is well beyond even the most optimistic expectations for next-generation centimeter-wavelength telescopes, such as the Phase 1 of the Square Kilometre Array (SKA1, cf. Cendes et al., 2022). Planets, or more likely proto-planets, likely also are to be detected at centimeter to millimeter wavelengths (e.g., Wolf, 2008; Pérez et al., 2016), with the the Very Large Array (VLA) or the Atacama Large Millimeter/submillimeter Array (ALMA), and there is increased potential of such studies with the next-generation Very Large Array (ngVLA). Clearly promising, but also poised for significant discoveries over the next few years, this chapter defers any discussion of radio emission at these wavelengths. Finally, this chapter does not consider the possibility of detecting artificial transmissions, though this approach may represent the ultimate in extrasolar planet discovery (Tarter, 2001).
@@ 0.2 Observational Constraints for Extrasolar Planets
  • Multiple observations have been conducted of M dwarfs and brown dwarfs, including of TRAPPIST-1, which have shown that they can produce coherent radio emission in the absence of a planet or that the presence of planets does not lead necessarily to magnetic star-planet interactions generating radio emissions (Pineda et al., 2017; Llama et al., 2018; Pineda & Hallinan, 2018; Villadsen & Hallinan, 2019; Callingham et al., 2021b).
@@ 0.5 Observational Considerations
Several observationally-relevant conclusions can be drawn, either from the heuristic approach described in the previous section or from a more exact treatment. First, ECMI planetary radio emission should be highly circularly polarized. This conclusion follows simply from the nature of cyclotron emission itself (producted by gyrations around magnetic field lines), and circularly polarized emission is the standard for planetary radio emissions in the Solar System. As few radio astronomical sources are circularly polarized, searching for circularly-polarized radio sources is a common technique for identifying candidate extrasolar planets (or stars) emitting via the ECMI (Callingham et al., 2023b).
@@ 0.6 Extensions to and Predictions for Extrasolar Planets
Finally, in the interest of completeness, planets in more “exotic” environments have been considered as possible radio emitters. These environments include planets around pulsars (Mishra et al., 2023), terrestrial planets around white dwarfs (Willes & Wu, 2005), planets around evolved cool stars (Ignace et al., 2010; Fujii et al., 2015), planets around T Tauri stars (Vidotto et al., 2010), and even interstellar “rogue planets,” i.e., planets not bound to a star (Vanhamäki, 2011). While there nature is not yet clear, Jupiter-Mass Binary Objects (JuMBOs) may be planets (Pearson & McCaughrean, 2023), and at least one such JuMBO has been detected at radio wavelengths (Rodriguez et al., 2024).
@@ 0.7 Planetary Magnetic Fields and Interiors
Emerging approachs to provide additional constraints on a planet’s interior structure are to assume that the composition of the planet is similar to that of its host star, at least for the refractory elements (Dorn et al., 2015, 2017; Brugger et al., 2017; Schulze et al., 2021; Unterborn et al., 2023), or can be inferred from atmospheric abundances (e.g., Bloot et al., 2023; Guimond et al., 2023). The technique of linking the host star-planet compositions likely requires high-precision spectroscopic measurements of the host star in order to obtain sufficient constraints on a planet’s composition (Wang et al., 2019). Moreover, there also are recent examples of planets that appear to have compositions discrepant, sometimes substantially so, from their host stars (e.g., Bean et al., 2023; Shi et al., 2023). There is even the possibility that a planet’s composition may reflect not only how far from its host star that it formed (radial location within the protoplanetary disk), but also azimuthal variations or other variations (e.g., temperature) within the protoplanetary disk (Keyte et al., 2023; van Dishoeck et al., 2023). Nonetheless, there also may be opportunities to model the likelihood of a planet generating a magnetic field using its host star metallicity as an additional constraint.
@@
Measurements of or constraints on the strengths of extrasolar planetary magnetic fields, even for a small number of extrasolar planets, would be valuable from two, complementary perspectives. First, the presence of a magnetic field requires the planet to support an internal dynamo, which in turn requires some internal region of the planet to support convection. Considering a planet with a given mass and radius, only a limited set of bulk compositions, thermal states, and internal pressure profiles may enable a dynamo region to persist (e.g., Yunsheng Tian & Stanley, 2013).
@@ 0.7.1 Gas-Giant Planets
For gas-giant planets (i.e., Jupiter mass), the equation of state of hydrogen is known sufficiently well to be confident that it undergoes a transition to a metallic state in a gas-giant planet’s interior (Hubbard et al., 2002; Helled et al., 2020). Further, Batygin (2018) argues that rotation rates of gas giants planets being well below their breakup velocities is consistent with magnetic braking of the planet due to its magnetic field being coupled to the (ionized) circumplanetary disk. Consequently, all gas-giant planets are expected to sustain planetary-scale magnetic fields, but there might be a wide range of magnetic field strengths. For instance, estimates for the polar magnetic field strength of HD 209458 b, achived with a variety of different (model-dependent) methods, have ranged over as much as two orders of magnitude (Sánchez-Lavega, 2004; Batygin & Stevenson, 2010; France et al., 2010; Ekenbäck et al., 2010; Khodachenko et al., 2021), from much smaller than that of Jupiter to much larger.
@@ 0.7.2 Ice-Giant Planets
A complementary approach is to obtain magnetic field measurements for even a modest sample of extrasolar ice giants. A testable hypothesis motivating such measurements could be that ice giants produce magnetic fields with a broad range of strengths, with Uranus and Neptune at the low end of this range. This broad range of magnetic field strengths could result from various origins or evolutions. For instance, ice giants forming at different distances from their host stars might have different bulk compositions, which could affect whether a water-rich or ammonia-rich internal layer generates the dynamo. Similarly, ice giants might produce a range of magnetic field topologies, ranging from the complex multi-polar topologies of Uranus and Neptune to more dipolar topologies characteristic of Jupiter and Earth. At a distance, dipolar toplogies would appear to have larger magnetic field strengths than multi-polar topologies.
@@ 0.7.3 Super Earths
Because the Solar System contains no sub-Neptunes nor super-Earths, there is an opportunity for discovery should magnetic fields be able to be measured for these classes of planets. Indeed, Coppari et al. (2021) have measured various properties of samples under the pressures and temperatures expected in the interiors of super-Earths. They find that the higher temperatures and pressures in super-Earths, relative to terrestrial-mass planets, could result in different mineralogic properties, including lower viscosities. A potential consequence, they speculate, could be that super-Earths would have an increased “electromagnetic coupling between the core and the mantle, enhancing convection and heat transport out of the core and affecting the strength and expression of the magnetic field.”
@@ 0.7.4 Terrestrial-Mass Planets
Noack & Lasbleis (2020) have shown that variations in the relative fractions of silicates and iron within terrestrial-mass planets could produce substantially different temperature-pressure profiles. While they do not extend their analysis to modeling the generation of a dynamo, differences in the interior state reasonably can be expected to result in different magnetic field strengths, topologies, or both. Indeed, estimates of the Earth’s paleomagnetic field show variations (in the median) of approximately a factor of a few over the past 3500 Myr (e.g., Bono et al., 2022), suggesting that the dynamo evolved (substantially) over gelogic time as its interior properties evolved. Further, Ziegler & Stegman (2013) and Blanc et al. (2020) illustrate how the Earth’s dynamo may have shifted from its mantle to its core as its interior cooled and the inner core began to solidify (Figure 5). If the dynamo has evolved over geological time scales, the Earth’s radio emissions may have had considerably different intensities as well.
@@ 0.8 Planetary Magnetic Fields and Habitability
There is a rich literature regarding how a planet’s magnetic field might protect its (secondary) atmosphere from erosion by its host star’s stellar wind or its surface from the effects of high-energy particles or both (Grießmeier et al., 2004; Grießmeier et al., 2005b; Lammer et al., 2008; Grießmeier et al., 2009; Grießmeier et al., 2010; Driscoll & Bercovici, 2013; Grießmeier et al., 2015; Foley & Driscoll, 2016; Garcia-Sage et al., 2017; Owen & Adams, 2019). The essential concept is that, if a planet’s magnetic field is sufficiently strong, and the kinetic and magnetic energies of the host star’s stellar wind are not too strong, the energetic particles in the host star’s stellar wind are deflected (by $\mathbf{v}\times\mathbf{B}$ forces) before they reach the planet’s atmosphere. This concept has been extended to consider the possibility that, even if a planet itself is not habitable, its magnetosphere could contribute to protecting a moon, thereby contributing to the moon’s habitability (Green et al., 2021).
@@
There have been efforts to assess the role of the Earth’s magnetosphere in determining its habitability (Varela et al., 2023) and speculation that changes in the Earth’s paleomagnetosphere may have contributed to substantial changes in the evolution of life (e.g., Meert et al., 2016). This expectation of atmospheric protection is has been extended to be included in models that attempt to assess or predict the likely habitability of extrasolar planets (Rodríguez-Mozos & Moya, 2017, 2019; McIntyre et al., 2019).

### v2

@@ Abstract
Detection of radio emission from Jupiter was identified quickly as being due to its planetary-scale magnetic field. Subsequent spacecraft investigations have revealed that many of the planets, and even some moons, either have or have had large-scale magnetic fields. In the case of the Earth, Jupiter, Saturn, Uranus, and Neptune, their magnetic fields are generated by dynamo processes within these planets, and an interaction between the solar wind and their magnetic fields generates intense radio emission via the electron cyclotron maser instability. In the case of Jupiter, its magnetic field interacts with the moon Io to result in radio emission as well. Extrasolar planets reasonably may be expected to generate large-scale magnetic fields and to sustain an electron cyclotron maser instability. Not only may these radio emissions be a means for discovering extrasolar planets, because magnetic fields are tied to the properties of planetary interiors, radio emissions may be a remote sensing means of constraining extrasolar planetary properties that will be otherwise difficult to access. In the case of terrestrial planets, the presence or absence of a magnetic field may be an indicator for habitability. Since the first edition of the Handbook, there have been a number of advances, albeit there remain no unambiguous detection of radio emission from extrasolar planets. New ground-based telescopes and new possibilities for space-based telescopes provide promise for the future.
@@ 0.1 Introduction
The focus of this chapter is on electron cyclotron maser instability (ECMI) emissions resulting from a stellar wind-planetary magnetosphere interaction. While planets, or at least Jupiter, can generate synchrotron emission that appears at decimeter and centimeter wavelengths, it is likely to remain well beyond detection capabilities in the near future. As an example, the centimeter-wavelength flux density of Jupiter is of order a few Janskys (e.g., Klein et al., 1989; de Pater et al., 2003). Scaled to a distance of even a few parsecs, the resulting flux density would be sub-nanoJansky, which is well beyond even the most optimistic expectations for next-generation centimeter-wavelength telescopes, such as the Phase 1 of the Square Kilometre Array (SKA1, cf. Cendes et al., 2022). Planets, or more likely proto-planets, likely also are to be detected at centimeter to millimeter wavelengths (e.g., Wolf, 2008; Pérez et al., 2016), with the the Very Large Array (VLA) or the Atacama Large Millimeter/submillimeter Array (ALMA), and there is increased potential of such studies with the next-generation Very Large Array (ngVLA). Clearly promising, but also poised for significant discoveries over the next few years, this chapter defers any discussion of radio emission at these wavelengths. Finally, this chapter does not consider the possibility of detecting artificial transmissions, though this approach may represent the ultimate in extrasolar planet discovery (Tarter, 2001).
@@ 0.2 Observational Constraints for Extrasolar Planets
  • Multiple observations have been conducted of M dwarfs and brown dwarfs, including of TRAPPIST-1, which have shown that they can produce coherent radio emission in the absence of a planet or that the presence of planets does not lead necessarily to magnetic star-planet interactions generating radio emissions (Pineda et al., 2017; Llama et al., 2018; Pineda & Hallinan, 2018; Villadsen & Hallinan, 2019; Callingham et al., 2021b, a).
@@ 0.5 Observational Considerations
Several observationally-relevant conclusions can be drawn, either from the heuristic approach described in the previous section or from a more exact treatment. First, ECMI planetary radio emission should be highly circularly polarized. This conclusion follows simply from the nature of cyclotron emission itself (produced by gyrations around magnetic field lines), and circularly polarized emission is the standard for planetary radio emissions in the Solar System. As few radio astronomical sources are circularly polarized, searching for circularly-polarized radio sources is a common technique for identifying candidate extrasolar planets (or stars) emitting via the ECMI (Callingham et al., 2023b).
@@ 0.6 Extensions to and Predictions for Extrasolar Planets
Finally, in the interest of completeness, planets in more “exotic” environments have been considered as possible radio emitters. These environments include planets around pulsars (Mishra et al., 2023), terrestrial planets around white dwarfs (Willes & Wu, 2005), planets around evolved cool stars (Ignace et al., 2010; Fujii et al., 2015), planets around T Tauri stars (Vidotto et al., 2010), and even interstellar “rogue planets,” i.e., planets not bound to a star (Vanhamäki, 2011). While there nature is not yet clear, Jupiter-Mass Binary Objects (JuMBOs) may be planets (Pearson & McCaughrean, 2023), and at least one such JuMBO has been detected at radio wavelengths (Rodríguez et al., 2024).
@@ 0.7 Planetary Magnetic Fields and Interiors
Emerging approaches to provide additional constraints on a planet’s interior structure are to assume that the composition of the planet is similar to that of its host star, at least for the refractory elements (Dorn et al., 2015, 2017; Brugger et al., 2017; Schulze et al., 2021; Unterborn et al., 2023), or can be inferred from atmospheric abundances (e.g., Bloot et al., 2023; Guimond et al., 2023). The technique of linking the host star-planet compositions likely requires high-precision spectroscopic measurements of the host star in order to obtain sufficient constraints on a planet’s composition (Wang et al., 2019). Moreover, there also are recent examples of planets that appear to have compositions discrepant, sometimes substantially so, from their host stars (e.g., Bean et al., 2023; Shi et al., 2023). There is even the possibility that a planet’s composition may reflect not only how far from its host star that it formed (radial location within the protoplanetary disk), but also azimuthal variations or other variations (e.g., temperature) within the protoplanetary disk (Keyte et al., 2023; van Dishoeck et al., 2023). Nonetheless, there also may be opportunities to model the likelihood of a planet generating a magnetic field using its host star metallicity as an additional constraint.
@@
Measurements of, or constraints on, the strengths of extrasolar planetary magnetic fields, even for a small number of extrasolar planets, would be valuable from two, complementary perspectives. First, the presence of a magnetic field requires the planet to support an internal dynamo, which in turn requires some internal region of the planet to support convection. Considering a planet with a given mass and radius, only a limited set of bulk compositions, thermal states, and internal pressure profiles may enable a dynamo region to persist (e.g., Yunsheng Tian & Stanley, 2013).
@@ 0.7.1 Gas-Giant Planets
For gas-giant planets (i.e., Jupiter mass), the equation of state of hydrogen is known sufficiently well to be confident that it undergoes a transition to a metallic state in a gas-giant planet’s interior (Hubbard et al., 2002; Helled et al., 2020). Further, Batygin (2018) argues that rotation rates of gas giants planets being well below their breakup velocities is consistent with magnetic braking of the planet due to its magnetic field being coupled to the (ionized) circumplanetary disk. Consequently, all gas-giant planets are expected to sustain planetary-scale magnetic fields, but there might be a wide range of magnetic field strengths. For instance, estimates for the polar magnetic field strength of HD 209458 b, achieved with a variety of different (model-dependent) methods, have ranged over as much as two orders of magnitude (Sánchez-Lavega, 2004; Batygin & Stevenson, 2010; France et al., 2010; Ekenbäck et al., 2010; Khodachenko et al., 2021), from much smaller than that of Jupiter to much larger.
@@ 0.7.2 Ice-Giant Planets
A complementary approach is to obtain magnetic field measurements for even a modest sample of extrasolar ice giants. A testable hypothesis motivating such measurements could be that ice giants produce magnetic fields with a broad range of strengths, with Uranus and Neptune at the low end of this range. This broad range of magnetic field strengths could result from various origins or evolutions. For instance, ice giants forming at different distances from their host stars might have different bulk compositions, which could affect whether a water-rich or ammonia-rich internal layer generates the dynamo. Similarly, ice giants might produce a range of magnetic field topologies, ranging from the complex multi-polar topologies of Uranus and Neptune to more dipolar topologies characteristic of Jupiter and Earth. At a distance, dipolar topologies would appear to have larger magnetic field strengths than multi-polar topologies.
@@ 0.7.3 Super Earths
Because the Solar System contains no sub-Neptunes nor super-Earths, there is an opportunity for discovery should magnetic fields be able to be measured for these classes of planets. Indeed, Coppari et al. (2021) have measured various properties of samples under the pressures and temperatures expected in the interiors of super-Earths. They find that the higher temperatures and pressures in super-Earths, relative to terrestrial-mass planets, could result in different mineralogical properties, including lower viscosities. A potential consequence, they speculate, could be that super-Earths would have an increased “electromagnetic coupling between the core and the mantle, enhancing convection and heat transport out of the core and affecting the strength and expression of the magnetic field.”
@@ 0.7.4 Terrestrial-Mass Planets
Noack & Lasbleis (2020) have shown that variations in the relative fractions of silicates and iron within terrestrial-mass planets could produce substantially different temperature-pressure profiles. While they do not extend their analysis to modeling the generation of a dynamo, differences in the interior state reasonably can be expected to result in different magnetic field strengths, topologies, or both. Indeed, estimates of the Earth’s paleomagnetic field show variations (in the median) of approximately a factor of a few over the past 3500 Myr (e.g., Bono et al., 2022), suggesting that the dynamo evolved (substantially) over geologic time as its interior properties evolved. Further, Ziegler & Stegman (2013) and Blanc et al. (2020) illustrate how the Earth’s dynamo may have shifted from its mantle to its core as its interior cooled and the inner core began to solidify (Figure 5). If the dynamo has evolved over geological time scales, the Earth’s radio emissions may have had considerably different intensities as well.
@@ 0.8 Planetary Magnetic Fields and Habitability
There is a rich literature regarding how a planet’s magnetic field might protect its (secondary) atmosphere from erosion by its host star’s stellar wind or its surface from the effects of high-energy particles or both (Grießmeier et al., 2004; Grießmeier et al., 2005b; Dehant et al., 2007; Lammer et al., 2008; Grießmeier et al., 2009; Grießmeier et al., 2010; Driscoll & Bercovici, 2013; Grießmeier et al., 2015; Foley & Driscoll, 2016; Garcia-Sage et al., 2017; Owen & Adams, 2019). The essential concept is that, if a planet’s magnetic field is sufficiently strong, and the kinetic and magnetic energies of the host star’s stellar wind are not too strong, the energetic particles in the host star’s stellar wind are deflected (by $\mathbf{v}\times\mathbf{B}$ forces) before they reach the planet’s atmosphere. This concept has been extended to consider the possibility that, even if a planet itself is not habitable, its magnetosphere could contribute to protecting a moon, thereby contributing to the moon’s habitability (Green et al., 2021).
@@
There have been efforts to assess the role of the Earth’s magnetosphere in determining its habitability (Varela et al., 2023) and even speculation that changes in the Earth’s paleomagnetosphere may have contributed to substantial changes in the evolution of life (e.g., Meert et al., 2016). This expectation of atmospheric protection is has been extended to be included in models that attempt to assess or predict the likely habitability of extrasolar planets (Rodríguez-Mozos & Moya, 2017, 2019; McIntyre et al., 2019).